You are given the following information: Stockholders' equity as reported on the firm's balance sheet = $4 billion, price/earnings ratio = 21.5, common shares outstanding = 24 million, and market/book ratio = 1.9. The firm's market value of total debt is $4 billion; the firm has cash and equivalents totaling $220 million; and the firm's EBITDA equals $1 billion. What is the price of a share of the company's common stock? Do not round intermediate calculations. Round your answer to the nearest cent. 


What is the firm's EV/EBITDA? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Stockholder's Equity as reported on the firm's Balance Sheet (Book Value) = $ 4 billion, Market/Book Ratio = 1.9

Market Value of Equity = 1.9 x 4 = $ 7.6 billion

Number of Shares Outstanding = 24 million

Price of Share of the firm's Common Stock = (7.6 x 1000) / 24 = $ 316.667 ~ $ 316.67

Market Value of Debt = $ 4 billion and Cash & Equivalents = $ 220 million or $ 0.22 billion

Enterprise Value = Market Value of Equity + Market Value of Debt - Cash & Equivalents = 7.6 + 4 - 0.22 = $ 11.38 billion

EBITDA = $ 1 billlion

Therefore, EV/EBITDA = 11.38 / 1 = 11.38